OpenStudy (anonymous):

What is the 17th term of the arithmetic sequence 18, 22, 26, …

OpenStudy (anonymous):

find the difference then plug it in to your formula

OpenStudy (anonymous):

COMMON DIFFERENCE, D = 4 AND IT IS AN A.P A = 18 A(17) = A+ 16 D = 18 +(16)4 = 18 + 64 = 82

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

yep!!!!

OpenStudy (anonymous):

thanks